# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

### v2

@@ Abstract
The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption. Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to runtime, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in. We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime. The VWT groups visual subwords (image patches) that are frequently used into visual words while infrequent ones remain intact. To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for sequence compression. Experimentally, we demonstrate a reduction in wattage of up to 25% with only a 20% increase in runtime at most. Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 100% or more). Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
@@
Figure 1: Visualization of patch matching by our visual word tokenizer (${\mathcal{T}}_{inter}^{100}$ model). Patches that are matched with one another are indicated by identical colors. Higher patch matching is exhibited by the background rather than foreground across datasets. Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
@@ 1 Introduction
In recent years, deep learning has seen continuous integration into a variety of systems worldwide. From coding to gaming, neural networks are increasingly deployed in online scenarios where asynchronous requests are processed in real-time. However, due to the size and complexity of modern architectures, such models are costly to run in practice. To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [24], Pruning [22, 41], and Quantization [11]. However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime. In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [18, 19, 9, 67, 43]. Newer large language models (LLMs) [58, 14] exhibit a noticeably larger vocabulary than their earlier counterparts [12, 45], thereby producing shorter sequences across various distributions. For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [13]. As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens. Similar to downsampling in convolutional neural networks, most research [27, 32, 48, 40, 5, 4, 26] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead. Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
@@
To introduce variability, we draw inspiration from subword tokenization algorithms [17, 51] used in NLP which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units. Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words. We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features while infrequent ones remain intact as they might carry task-relevant information. We propose two procedures to capture this principle. The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g. backgrounds). The second is an inter-image approach where basic features across multiple images such as colors or edges are discovered as visual words. Image patches are then grouped based on the similarity of these basic characteristics. Crucially, patches that have distinct characteristics (i.e. high dissimilarity with any visual word) remain intact and form separate visual subwords. The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
@@ 2 Related Work
Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores. In Liang et al. [32], Xu et al. [66], Kong et al. [27], Bian et al. [4], redundancy is addressed by fusing tokens. Both Rao et al. [48] and Tang et al. [55] opt to prune such tokens instead. Recent efforts [7, 8, 6, 26] attempt to combine the benefits of merging and pruning. In Tran et al. [60], an additional metric termed the energy score is used to better identify redundancy. Uniquely, Fayyaz et al. [16] use inverse transform sampling to select important tokens. Most relevant to our work are Marin et al. [40] and Bolya et al. [5]. The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike Bolya et al. [5], we do not include Marin et al. [40] as one of our baselines due to a lack of code release..
Figure 2: Overview of the Visual Word Tokenizer. An intra-image approach. In the forward pass, the pixel variance of the patches is computed with the top-k lowest values being masked (deploy). The grouped tokens are dropped after positional information is added. Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used. An inter-image approach. A Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process). In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy). The grouped tokens are averaged after positional information is added.
@@
Our method also takes inspiration from efficient inference in NLP. Increasingly, the tokenizer’s vocabulary is specialized to reduce the input token length. In Gee et al. [18], domain adaptation of the tokenizer ensures fewer subword or character tokens are produced. Gee et al. [19] followed up by introducing n-grams for tokenization beyond the word-level boundary. In Dagan et al. [9], tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs. Meanwhile, Yamaguchi et al. [67] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference. Recently, Minixhofer et al. [43] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
Dataset Model $Base$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@
Table 1: Token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. Unlike text tokenizers, domain specialization (i.e. In-Domain) does not result in greater compression. Smaller vocabularies produce shorter sequences whereas larger vocabularies result in patches being increasingly matched to separate visual words.
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
Waterbirds CLIP 123.99 73.89 91.64 110.93 102.65 107.93 117.36
CelebA 126.53 74.47 93.07 114.37 96.31 102.12 115.46
MetaShift 123.30 74.02 90.96 113.84 98.98 105.78 114.70
OpenImages (Com.) 135.05 74.80 94.63 114.62 106.33 112.67 123.95
OpenImages (Rare) 135.66 75.27 94.04 115.78 104.73 110.75 123.05
COCO BLIP 191.74 81.68 147.23 169.77 144.71 156.25 171.84
NoCaps 185.12 81.89 149.26 175.06 141.20 155.11 170.38
(a) Wattage (watt) $\downarrow$
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
Waterbirds CLIP 8.48 88.55 17.04 5.84 9.88 9.73 9.91
CelebA 8.33 89.41 17.25 5.86 9.45 9.54 9.99
MetaShift 8.21 89.48 17.23 6.01 9.99 9.87 9.75
OpenImages (Com.) 8.83 92.23 17.44 5.97 9.64 9.69 10.23
OpenImages (Rare) 8.60 86.00 17.04 6.19 9.69 10.02 9.67
COCO BLIP 12.45 57.11 15.06 8.10 8.73 9.10 10.31
NoCaps 12.44 57.61 15.08 8.04 9.49 9.28 10.51
(b) Runtime (millisecond) $\downarrow$
Table 2: Wattage and runtime per sample. Compared to $Base$, VWTs reduce wattages by up to 25% with ${\mathcal{T}}_{inter}^{100}$. Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former. ${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{10000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
Vector Quantization.
The idea of discretizing continuous distributions has been explored in many works, most recently for image generation. Yang et al. [68] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words. In Wu et al. [64] and Bao et al. [3], discretization is applied as part of the modelling for ViTs. Van Den Oord et al. [61] learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach. Esser et al. [15] and Yu et al. [71] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework. Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [47, 72, 38, 39, 57, 56, 54]. Lastly, Yang et al. [69] tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
3 Visual Word Tokenizer
@@
In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings. The number of patches is typically fixed (e.g. $197$) based on the choice of architecture. We seek to split an image into variable length inputs instead (i.e. certain images will use $80$ tokens, while others a $100$). This variability is induced by conventional text tokenizers [17, 51, 65, 28] for model efficiency. We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the pixel space. These visual words capture frequently used patches while retaining infrequent ones as is. A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach). Figure 2 summarizes the Visual Word Tokenizer (VWT).
@@
Table 3: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP). VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds. On COCO and NoCaps, VWTs retain a higher performance than $ToME$.
@@ 3.1 An intra-image approach
The pixel variance of the image patches is the simplest criterion that can be used for grouping. In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest intact. To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity. We do not include the [CLS] token for dropping. This approach is inspired by Minderer et al. [42] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector. Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thereby increasing training efficiency.
@@ 3.2 An inter-image approach
Inspired by text tokenizers and codebooks, we propose a variable-length approach that statistically discovers visual words across many images. The tokenizer consists of two steps: Pre-Process and Deploy
Figure 3: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model). Patches with the lowest pixel variance that are dropped are indicated in black. In most cases, the dropped patches correspond to the background object on average which is typically uninformative.
@@
The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations. In Yang et al. [68], k-means clustering is applied to keypoint descriptors from SIFT [37] to learn a fixed set of centroids. These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ). In our method, we adopt a variation of this framework by building the BoVW using patches within the pixel space. Our design choice is motivated by two main factors. First, we find keypoint descriptors to be costly for inference. In each forward pass, computing keypoints for each image would significantly increase runtime. Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another. Such issues were also described by Bolya et al. [5], thus leading to their use of attention scores instead. Further justification for leveraging the pixel space is provided in Section 4.6.
Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g. 16 for ViT-B/16). We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word. Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings. We also execute this process in batches using the MiniBatchKMeans333from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints. Note that MiniBatchKMeans uses the Euclidean distance by design. Since clustering is done in the pixel space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
@@
For the inter-image approach, if batching instead of online inference is desired, the uniform requirement of tensors becomes a challenge. To maintain parallelism, any reduction has to be equal across samples. Due to the non-uniformity of tokenization, sequences have to be sequentially compressed before padding to the same length. We opt to append additional [PAD] tokens until the longest length within the batch is achieved. Similar to text transformers [62], the attention scores are set to negative infinity before the softmax to nullify the influence of padding. We do not add positional information to the [PAD] tokens as extrapolating such information to non-uniform sequences will significantly worsen model efficiency.

### v4

# Visual-Word Tokenizer: Beyond Fixed Sets of Tokens in Vision Transformers

@@ Abstract
The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption. Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to energy efficiency, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in. We introduce the Visual-Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance. The VWT groups visual subwords (image patches) that are frequently used into visual words, while infrequent ones remain intact. To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for sequence compression. Experimentally, we demonstrate a reduction in energy consumed of up to 47%. Comparative approaches of 8-bit quantization and token merging can lead to significantly increased energy costs (up to 500% or more). Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance. The experimental code for our paper is also made publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
@@
Figure 1: Visualization of the inter-image approach by the Visual-Word Tokenizer (${\mathcal{T}}_{inter}^{100}$ model). Patches that are matched with one another are indicated by identical colors. Higher patch matching is exhibited with the background rather than foreground object across the datasets. Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts during inference.
@@ 1 Introduction
In recent years, deep learning has seen continuous integration into a variety of systems worldwide. From coding to gaming, neural networks are increasingly deployed in online scenarios where asynchronous requests are processed in real-time. However, due to the size and complexity of modern architectures, such models are costly to run in practice. To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [hinton2015distilling], Pruning [han2015learning, michel2019sixteen], and Quantization [dettmers2022gpt3]. Many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce energy efficiency. In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [gee2022fast, gee2023multi, dagan2024getting, yamaguchi2024empirical, minixhofer2024zero]. Newer large language models (LLMs) [team2024gemma, dubey2024llama] exhibit a noticeably larger vocabulary than their earlier counterparts [devlin2018bert, radford2019language], thereby producing shorter sequences across various distributions. For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [dosovitskiy2020image]. As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens. Similar to downsampling in convolutional neural networks, most research [kong2022spvit, liang2022not, rao2021dynamicvit, marin2021token, bolya2022token, bian2023multi, kim2024token] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead. Given the analogous architecture of the transformer across modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
@@
To introduce variability, we draw inspiration from subword tokenization algorithms [gage1994new, sennrich2016neural] used in NLP, which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units. Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words. We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features, while infrequent ones remain intact as they might carry task-relevant information. We propose two procedures to capture this principle. The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g., backgrounds). The second is an inter-image approach where basic features across multiple images such as colors or edges are discovered as visual words. Image patches are then grouped based on the similarity of these basic characteristics. Crucially, patches that have distinct characteristics (i.e., high dissimilarity with any visual word) remain intact and form separate visual subwords.
@@ 2 Related Work
Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores. In liang2022not, xu2022evo, kong2022spvit, bian2023multi, redundancy is addressed by fusing tokens. Both rao2021dynamicvit and tang2022patch opt to prune such tokens instead. Recent efforts [cao2023pumer, chen2023diffrate, bonnaerens2023learned, kim2024token] attempt to combine the benefits of merging and pruning. In tran2024accelerating, an additional metric termed the energy score is used to better identify redundancy. Uniquely, fayyaz2022adaptive use inverse transform sampling to select important tokens. Most relevant to our work are marin2021token and bolya2022token. The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike bolya2022token, we do not include marin2021token as one of our baselines due to a lack of code release..
Figure 2: Overview of the Visual-Word Tokenizer (VWT). An intra-image approach (deploy only). During inference, the pixel variance of the patches is computed with the top-k lowest values being masked. The masked tokens are dropped after positional information is added. An inter-image approach (pre-process & deploy). First, a Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process). Then, during inference, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy). The unmasked tokens are averaged based on their grouping to the same visual word after positional information is added.
@@
Our method also takes inspiration from efficient inference in NLP. Increasingly, the tokenizer’s vocabulary is specialized to reduce the input token length. In gee2022fast, domain adaptation of the tokenizer ensures fewer subword or character tokens are produced. gee2023multi followed up by introducing n-grams for tokenization beyond the word-level boundary. In dagan2024getting, tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs. Meanwhile, yamaguchi2024empirical analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference. Recently, minixhofer2024zero leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
Vector Quantization.
The idea of discretizing continuous distributions has been explored in many works, most recently for image generation. yang2007evaluating leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words. In wu2020visual and bao2021beit, discretization is applied as part of the modelling for ViTs. van2017neural learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach. esser2021taming and yu2021vector further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework. Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [ramesh2021zero, yu2022scaling, lu2022unified, lu2024unified, team2023gemini, team2024chameleon, sun2024autoregressive]. Lastly, yang2022visual tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
3 Visual-Word Tokenizer
Dataset Model BaseBase ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@ 3 Visual-Word Tokenizer
Table 1: Token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. VWTs significantly reduce the token length of $Base$, particularly for larger image sizes (384 $\times$ 384 on COCO and NoCaps). Unlike text tokenizers, domain specialization (In-Domain) does not result in greater compression for the interinter-image approach. Smaller vocabularies produce shorter sequences whereas larger vocabularies result in patches being increasingly matched to separate visual words.
Dataset Model BaseBase $Q_{8}$ ToMeToMe ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
Waterbirds CLIP 1.05 6.54 1.56 0.65 1.01 1.05 1.16
CelebA 1.05 6.66 1.61 0.67 0.91 0.97 1.15
MetaShift 1.01 6.62 1.57 0.68 0.99 1.04 1.12
OpenImages (Com.) 1.19 6.90 1.65 0.68 1.03 1.09 1.27
OpenImages (Rare) 1.17 6.47 1.60 0.72 1.01 1.11 1.19
COCO BLIP 2.39 4.66 2.22 1.38 1.26 1.42 1.77
NoCaps 2.30 4.72 2.25 1.41 1.34 1.44 1.79
Table 2: Energy (joule) consumed per sample. The power and runtime values that were used to compute the energy consumed are provided in Table 9. Compared to $Base$, VWTs reduce energy usage by up to 43% with ${\mathcal{T}}_{intra}^{0.5}$ and 47% with ${\mathcal{T}}_{inter}^{100}$. Both $Q_{8}$ and ToMeToMe significantly increase energy costs, particularly on datasets with smaller image sizes (224 $\times$ 224 on Waterbirds, CelebA, MetaShift, OpenImages). For the inter-image approach, efficiency improvements are highest with larger image sizes (384 $\times$ 384 on COCO and NoCaps) and decreases naturally with larger vocabularies due to smaller compression.
@@
In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings. The number of patches is typically fixed (e.g., $197$) based on the choice of architecture. We seek to split an image into variable length inputs instead (i.e., certain images will use $80$ tokens, while others a $100$). This variability is induced by conventional text tokenizers [gage1994new, sennrich2016neural, wu2016google, kudo2018sentencepiece] for model efficiency. We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the pixel space. These visual words capture frequently used patches while retaining infrequent ones as is. A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach). Figure 2 summarizes the Visual-Word Tokenizer (VWT).
@@
Table 3: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP). $Q_{8}$ displays the lowest degradation in performance relative to $Base$, but suffers from a significant increase in energy used as shown in Table 2. The VWTs are competitive with $ToMe$, particularly for image captioning on COCO and NoCaps. Both $Q_{8}$ and the VWTs are shown to improve the worst-group accuracies on Waterbirds and CelebA. For the latter, the intraintra-image approach works best on MetaShift and OpenImages while the interinter-image approach is more suitable on the remaining datasets. The lower performance of the interinter-image approach on MetaShift and OpenImages may stem from the lack of uniform backgrounds (e.g., skies in Waterbirds) that increasingly shifts compression to the foreground object.
@@ 3.1 An intra-image approach
The pixel variance of the image patches is the simplest criterion that can be used for grouping. In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest intact. To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity. We do not include the [CLS] token for dropping. This approach is inspired by minderer2024scaling which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector. Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thereby increasing training efficiency.
@@ 3.2 An inter-image approach
Inspired by text tokenizers and codebooks, we propose a variable-length approach that statistically discovers visual words across many images. The tokenizer consists of two steps: Pre-Process and Deploy.
Figure 3: Visualization of the intra-image approach by the Visual-Word Tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model). Patches with the lowest pixel variance that are dropped are indicated in black. In most cases, the dropped patches correspond to the uniform background which is uninformative. However, in a few cases, the patches of the foreground object are dropped (e.g., airplane in first image for COCO) which is undesirable.
@@
The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations. In yang2007evaluating, k-means clustering is applied to keypoint descriptors from SIFT [lowe2004distinctive] to learn a fixed set of centroids. These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ). In our method, we adopt a variation of this framework by building the BoVW using patches within the pixel space. Our design choice is motivated by two main factors. First, we find keypoint descriptors to be costly for inference. In each forward pass, computing keypoints for each image would significantly increase runtime. Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another. Such issues were also described by bolya2022token, thus leading to their use of attention scores instead. Further justification for leveraging the pixel space is provided in Section 4.4.
Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g., 16 for ViT-B/16). We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word. Patchification is done via basic tensor operations [^3] We decompose the image into smaller tensors given the patch size while avoiding any linear transformations. and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings. We also execute this process in batches using the MiniBatchKMeans444from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints. Note that MiniBatchKMeans uses the Euclidean distance by design. Since clustering is done in the pixel space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
Figure 5: Visualization of long-form captioning on COCO. Longer captions are generated via a length penalty of 2.0 and a maximum length of 40. Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its higher compression.
@@
For the inter-image approach, if batching instead of online inference is desired, the uniform requirement of tensors becomes a challenge. To maintain parallelism, any reduction has to be equal across samples. Due to the non-uniformity of tokenization, sequences have to be sequentially compressed before padding to the same length. We opt to append additional [PAD] tokens until the longest length within the batch is achieved. Similar to text transformers [vaswani2017attention], the attention scores are set to negative infinity before the softmax to nullify the influence of padding. We do not add positional information to the [PAD] tokens as extrapolating such information to non-uniform sequences will significantly worsen model efficiency.